Fermi National Laboratory


Accelerator Update

Wednesday July 28
The day shift began with Operations monitoring store 3678 and the Recycler stash, and with the Antiproton Source (Pbar), Booster, Switchyard (SY), Linac, and MiniBooNE in access for maintenance work.

Linac came out of access around 9:50 AM.

The Neutron Therapy Facility (NTF) began a study period at 10:23 AM.

Booster came out of access at 10:54 AM.

Operations established beam to the Main Injector dump at 12:30 PM.

Pbar began stacking at 1:05 PM.

Operations established beam to the Meson MCenter experiment at 1:09 PM.

Operations established beam to the Meson MTest experiment at 3 PM.

Cryo system technicians at CHL reported at 9:20 PM that they were hit by a power glitch. The Duty Electrician informed the high voltage tech.

Thursday July 29
The midnight shift began with Operations monitoring store 3678 and the stash, with Pbar stacking, with MiniBooNE pending Horn repairs, and with MTest and MCenter taking beam.

Operations reported at 1:14 AM that they were having problems with MECAR.

An MTest experimenter said that they were done with the beam at 1:22 AM.

At 1:41 AM, the MI abort dropped on the MECAR bit. Operations tried a reboot, but it wouldn't take. Operators called in an expert who arrived and began his investigation around 2 AM.

Operations reported at 4:11 AM that the MECAR expert got it running, but when Operations tried to run MI beam there was still problems.

At 4:23 AM, just to complicate things, Operations discovered that a quadrupole power supply (MIQ10) was putting out more voltage than required. Operators called in a power supply expert.

At 4:29 AM, Operations noted that the Recycler stash was being affected by a MI Time Line Generator (TLG) problem.

A power supply expert reported at 6 AM that he replaced a cable driver card on MIQ10.

At 6:05 AM, Operations reported a problem with the MI beam. Operators called a MI expert who believed it was a RF problem. Due to this MI RF problem, there was no stacking or SY beam.

Two MI experts reported at 9:20 AM that they found a power supply problem and that a MECAR tune table had corrupted data. They restored the table and MI resumed normal beam operations.

Pbar resumed stacking at 9:24 AM.

Tevatron (TeV) experts began an end-of-store study around noon. The study ended at 1 PM and Operations terminated store 3678.

Operations reestablished beam to MCenter at 1:09 PM.

Operators had the Collider Detector at Fermilab (CDF) and D-Zero (D0) ready for an access by 1:16 PM.

Operations established beam to MTest at 1:35 PM.

At 2:02 PM, a Booster expert changed a regulation card for a power supply (GMPS). Booster was back up twenty minutes later.

Both CDF and D0 were out of access by 3:14 PM.

At 3:55 PM, as Operations was just starting a TeV dry squeeze, a cryo magnet (B0Q6) quenched.

MI studies ended at 5:55 PM.

Operations put the TeV through another dry squeeze at 7:31 PM and then through a wet squeeze at 8:24 PM.

Operations put the TeV back into shot setup at 8:54 PM.

At 10:46 PM, a Booster power supply (WAPS) tripped off. Operators investigated.

Operations established store 3685 at 11:22 PM with an initial luminosity of 71.8E30. Stacking could not resume due to Booster power supply problems.

Booster experts reported at 11:38 PM that the WAPS problem was due to a 480V West Gallery Yard breaker trip.

Friday July 30
The midnight shift began with Operations monitoring store 3685 and the stash, but with no beam beyond the Booster.

Operations reported at 12:06 AM that they had contacted the Duty Electrician and the High Voltage Coordinator. In addition, RF and power supply experts were coming in to help with the 480V turn on.

Pbar resumed stacking at 12:44 AM.

Operations reported at 5:28 AM that they finished stashing antiprotons into the Recycler, and that Pbar had resumed stacking and the MCenter experiment was taking beam.

MTest requested beam at 6:50 AM.

Plans for Friday and the Weekend
The Run Coordinator says the store will remain in the TeV until early on the Saturday day shift. A Linac RF station (LRF5) needs its power amplifier changed; this will be done this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/30/2004   email Fermilab